OpenStudy (marcelie):

help please !!!! For the following exercises, find the x- or t-intercepts of the polynomial functions. help please !!! For the following exercises, find the x- or t-intercepts of the polynomial functions. f (x) = x6 − 2x4 − 3x2

OpenStudy (anonymous):

To find the intercepts of the function f(x), then you have to put in f(0), and solve it. And to find the intercepts of x, then you have to put x6-2x4-3x2 = 0 and solve x

OpenStudy (marcelie):

then

zepdrix (zepdrix):

\[\large\rm f(x)=x^6−2x^4−3x^2\]So to find your x-intercepts, this is when our function intercepts/crosses the x-axis. This happens, as Tommy indicated, when y=0.\[\large\rm 0=x^6−2x^4−3x^2\]Solve for x! \c:/

zepdrix (zepdrix):

Start by pulling the greatest common factor out of each term. They all share something in common.

OpenStudy (marcelie):

can u please help me with multiplicty